How would you prepare a 1 M solution of ethanol from a 90% (v/v) stock? (MW = 46, d = 0.789)
Answer #1

90% v/v means that solution has 90 ml of ethanol and 100 ml of water

For Stock solution : Molarity = (m*1000)/ [M * vol (in ml)]

{ M = Molar mass of ethanol, vol = vol. of solution, m = given mass of ethanol}

also m = density * volume

=> Molarity = (0.789 * 90 * 1000) / (46 * 190)

=> Molarity = 8.124 mol/L

Let the solution we want to prepare is 250 ml

then, M1V1 = M2V2

{1 = solution to be prepared

2 = stock solution}

1 * 250 = 8.124 * x     [ x = vol. of stock solution required ]

=> x = 30.77

x = 30.7 ml of the stock solution then make up the volume to 250 ml by adding 219.3 ml of water.